OpenStudy (anonymous):

What is the slope of the line? 4x-5y=20

OpenStudy (imstuck):

Isolate the y term. Get it all alone on the left side of the equals sign and the coefficient in front of the x after all that is done is the slope of the line. It is NOT 4, just so you know; it's not that easy! I wish!

OpenStudy (anonymous):

-5y=-4x+20 y=(4/5)x-4 slope =4/5